# How Uniform Random Weights Induce Non-uniform Bias: Typical Interpolating Neural Networks Generalize with Narrow Teachers

### v2

@@ B.4 Extension to Non-Interpolators
Let $\varepsilon>0$. Assume that there exists a narrow teacher NN $h^{\star}$ s.t $\mathcal{L_{D}}\left(h^{\star}\right)=\varepsilon^{\star}>0$. Denote by $\tilde{p}$ the teacher equivalence probability, and define the posterior distribution
$\mathcal{P_{S}}(h)=\mathcal{P}_{\mathcal{S}}^{\varepsilon}(h)=\mathbb{P}\left(h\,\middle|\,\mathcal{L_{S}}(h)\leq\varepsilon^{\star}+\varepsilon\right)\,.$
In the G&C formulation, sampling from $\mathcal{P}_{\mathcal{S}}^{\varepsilon}$ is equivalent to stopping at the first model satisfying $\mathcal{L_{S}}\left(h\right)\leq\varepsilon^{\star}+\varepsilon$.
@@
Part 1: We will show that
@@
Since $h^{\star}\in\mathcal{H}_{\tau}$, w.p. $1-e^{-2\varepsilon^{2}N}$ it holds that $\mathcal{L_{S}}\left(h^{\star}\right)\leq\varepsilon^{\star}+\varepsilon$ so there exists such $\tilde{h}\in\mathcal{H}_{\tau}$. Then
@@
Notice that the sample complexity is quadratically dependent on the population error, as opposed to the linear dependence in the realizable case, i.e. when the teacher is assumed to be a perfect interpolator.
@@ B.5 Relationship to PAC Bayes
Discussion. Theorem B.15 guarantees generalization with probability at least $1-\delta_{h}$ over the hypothesis sampling and probability $\delta_{\mathcal{S}}$ over the data sampling. This separation between $\delta_{h}$ and $\delta_{\mathcal{S}}$ highlights how both sources of randomness play a role in generalization. Interestingly, the sample complexity term $N$ exhibits a logarithmic dependence on $\delta_{\mathcal{S}}$ and only a doubly logarithmic dependence on $\delta_{h}$. Thus, for any $\delta_{\mathcal{S}}$ and $\varepsilon$, the probability of not sampling a PAC interpolator decays extremely fast (doubly exponential in $N$):
@@ B.7 Proofs Using Nonuniform Learnability
In the next pages, we will show a result that does not use the teacher assumption for the generalization of randomly sampled networks. The result is more general than Lemma 3.2 and B.15, which were both tailored for the teacher assumption. However, the price to pay for relaxing this assumption is that the following result is slightly less tight. Since we do not use the teacher assumption, with some abuse of notation we use $\mathcal{D}$ to denote the joint distribution of feature-label pairs $\left(\mathbf{x},y\right)$.
@@
Proof.
[[[[of Theorem B.18]]]] From Lemma B.20 we have that
@@ D.2 Setting and Notation
The following generalization bound follows directly from Prop. B.23 and Theorem 5.6.
@@
We denote this event by $\mathcal{E}_{1}$. Recalling that $\hat{C}^{\mathrm{cont}}\geq-\log\left(\hat{p}_{\mathcal{S}}\right)$ when conditioned on $\mathcal{E}_{1}$, from Prop. B.23 we deduce that,

### v3

@@ B.4 Extension to Non-Interpolators
In this section we extend our results to the case where both the teacher and student are not assumed to be interpolating.
Specifically, we are interested in the following setting. Assume that there exists a narrow teacher NN $h^{\star}$ s.t ${\mathcal{L_{D}}\left(h^{\star}\right)=\varepsilon^{\star}>0}$. Denote by $\tilde{p}$ the teacher equivalence probability, and, given $\gamma>0$, define the posterior distribution
$\mathcal{P_{S}}(h)=\mathcal{P}_{\mathcal{S}}^{\gamma}(h)=\mathbb{P}\left(h\,\middle|\,\mathcal{L_{S}}(h)\leq\gamma\right)\,.$
In the G&C formulation, sampling from $\mathcal{P}_{\mathcal{S}}^{\gamma}$ is equivalent to stopping at the first model satisfying $\mathcal{L_{S}}\left(h\right)\leq\gamma$.
Theorem B.9.
Assume that there exists a teacher $h^{\star}$ with $\mathcal{L}_{\mathcal{D}}(h^{\star})=\varepsilon^{\star}>0$ and TE probability $\tilde{p}\in\left(0,\frac{1}{2}\right)$. Let $\varepsilon\in\left(0,\frac{1}{2}-\varepsilon^{\star}\right)$, $\delta\in\left(0,\frac{1}{5}\right)$, and $\gamma=\varepsilon^{\star}+\varepsilon$. If
$N\geq\frac{-\log\left(\tilde{p}\right)+3\log\left(\frac{2}{\delta}\right)}{2\varepsilon^{2}}$
then
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}^{\gamma}}\left(\mathcal{L_{D}}\left(h\right)\leq\gamma+\varepsilon\right)\geq 1-\delta\,.$
We proceed to state Hoeffding’s inequality, which is used in the proof of Theorem B.9, then continue to prove some lemmas leading to the theorem.
@@
The proof comprises standard arguments from the uniform convergence theory of generalization.
@@
Part 1: We show that
@@
Using Hoeffding’s inequality, it holds w.p. $1-e^{-2\varepsilon^{2}N}$ that $\mathcal{L_{S}}\left(h^{\star}\right)\leq\varepsilon^{\star}+\varepsilon$, and since $h^{\star}\in\mathcal{H}_{\tau}$, there exists $\tilde{h}\in\mathcal{H}_{\tau}$ satisfying the condition. Then
@@
Notice that the sample complexity is quadratically dependent on the population error $\varepsilon$, as opposed to the linear dependence in the realizable case, i.e. when the teacher is assumed to be a perfect interpolator. This is expected as this there is no realizability assumption.
@@ B.5 Relationship to PAC Bayes
Discussion.
Theorem B.16 guarantees generalization with probability at least $1-\delta_{h}$ over the hypothesis sampling and probability $\delta_{\mathcal{S}}$ over the data sampling. This separation between $\delta_{h}$ and $\delta_{\mathcal{S}}$ highlights how both sources of randomness play a role in generalization. Interestingly, the sample complexity term $N$ exhibits a logarithmic dependence on $\delta_{\mathcal{S}}$ and only a doubly logarithmic dependence on $\delta_{h}$. Thus, for any $\delta_{\mathcal{S}}$ and $\varepsilon$, the probability of not sampling a PAC interpolator decays extremely fast (doubly exponential in $N$):
@@ B.7 Proofs Using Nonuniform Learnability
In the next pages, we will show a result that does not use the teacher assumption for the generalization of randomly sampled networks. The result is more general than Lemma 3.2 and B.16, which were both tailored for the teacher assumption. However, the price to pay for relaxing this assumption is that the following result is slightly less tight. Since we do not use the teacher assumption, with some abuse of notation we use $\mathcal{D}$ to denote the joint distribution of feature-label pairs $\left(\mathbf{x},y\right)$.
@@
Proof of Theorem B.19.
From Lemma B.21 we have that
@@ D.2 Setting and Notation
The following generalization bound follows directly from Prop. B.24 and Theorem 5.6.
@@
We denote this event by $\mathcal{E}_{1}$. Recalling that $\hat{C}^{\mathrm{cont}}\geq-\log\left(\hat{p}_{\mathcal{S}}\right)$ when conditioned on $\mathcal{E}_{1}$, from Prop. B.24 we deduce that,